OpenStudy (anonymous):

What is the 12th term of the sequence 10, 6, 2, …? I got -34 am I right

jimthompson5910 (jim_thompson5910):

-34 is correct

jimthompson5910 (jim_thompson5910):

an = a1 + d(n-1) a12 = a1 + d(12-1) a12 = 10 + (-4)(12-1) a12 = -34